# QuickQuakeBuildings: Post-earthquake SAR-Optical Dataset for Quick Damaged-building Detection

### 11 Dec 2023 v1

@@ Abstract
This is the preprint version. Quick and automated earthquake-damaged building detection from post-event satellite imagery is crucial, yet it is challenging due to the scarcity of training data required for developing robust algorithms. This letter presents the first dataset dedicated to detecting earthquake-damaged buildings from post-event very high resolution (VHR) Synthetic Aperture Radar (SAR) and optical imagery. Utilizing open satellite imagery and annotations acquired after the 2023 Turkey–Syria earthquakes, we deliver a dataset of co-registered building footprints and satellite image patches of both SAR and optical data, encompassing more than four thousand buildings. The task of damaged building detection is formulated as a binary image classification problem, that can also be treated as an anomaly detection problem due to extreme class imbalance. We provide baseline methods and results to serve as references for comparison. Researchers can utilize this dataset to expedite algorithm development, facilitating the rapid detection of damaged buildings in response to future events. The dataset and codes together with detailed explanations are made publicly available at https://github.com/ya0-sun/PostEQ-SARopt-BuildingDamage.
@@ I Introduction
Earthquakes can result in substantial structural and infrastructural damage, often with significant socioeconomic consequences. After an event, fast and accurate detection of earthquake-damaged buildings in remote sensing imagery is of great importance. Remote sensing technologies can effectively improve the efficiency of disaster management and have been employed to estimate the extent of earthquake damage to buildings [1, 2, 3]. Very high resolution (VHR) optical images are easier to interpret, making them a preferred choice for many studies, if available. SAR imagery, on the other hand, is particularly interesting due to its capability of acquiring information regardless of weather and sun illumination conditions. Moreover, the enhanced resolution of contemporary SAR satellite images enables the extraction of information at the individual building level [4, 5, 6], comparable to VHR optical data.
Over the past years, many researchers have developed algorithms utilizing SAR data to detect earthquake-damaged buildings. Most existing works detect changes using both pre- and post-earthquake SAR images to acquire building damage information [7, 8]; however, pre-event high-resolution SAR imagery is generally unavailable in most locations. A few works simulate a SAR image using building shapes extracted from a pre-event optical image and acquisition parameters of a post-event SAR image and detect earthquake-damaged buildings by comparing the simulated and real SAR images  [9]. Utilizing only a single post-event SAR image, some researchers detect damaged buildings by analyzing signatures of destroyed buildings in high-resolution SAR data [10, 11, 12]. However, the study areas often comprise a limited number of isolated buildings, which do not accurately depict the typical conditions in densely populated urban areas. Therefore, the question of whether a single VHR SAR image acquired after an event can effectively identify damaged buildings remains to be addressed. In this regard, benchmark datasets play a pivotal role in the development and comparative assessment of diverse methodologies aimed at addressing the following questions: To what extent can a single VHR SAR image, acquired post-event, allow to identify damaged buildings and with which accuracy? Additionally, how do the outcomes derived from a single post-event SAR image compare with those of an optical image?
@@ II-A Study Area and Data Sources
The study area is chosen in the city of Islahiye, located in southeastern Turkey near the northwestern border of Syria. On February 6, 2023, a devastating magnitude 7.8 earthquake struck the Turkish province of Kahramanmaras, followed by a 7.5 magnitude aftershock nine hours later. The earthquake inflicted widespread destruction, resulting in a significant loss of life, numerous injuries, and extensive damage to buildings and infrastructure. Islahiye was one of the most affected areas.
@@ II-B2 Registering corresponding features
Since the SAR image we used, i.e., the GEO product, has been terrain corrected, we neglect global registration and perform subarea and polygon registration. The direction of transformation, GIS to SAR or SAR to GIS, is inconsequential for later analysis, and thus, we add the transformations of each step to building polygons. Afterward, manual validation and editing with expert knowledge are conducted to ensure accurate registration within the study area. Figure 2 shows building polygons on the SAR image in exemplary areas before and after registration. As can be seen, the registration procedure effectively aligns the two data sets. Note that buildings located around the mountainous area on the west side of the city, e.g., the area in and near the red box shown in Figure 1, are discarded due to unclear signals in the SAR image, which makes it difficult to check registration results.
@@ II-C Patch generation
Among the 4029 buildings in the study area, 169 are damaged, and the remaining 3860 buildings are intact. Consequently, the dataset contains 169 damaged buildings and 3860 intact buildings, and each of them has four patches: a SAR image patch, a SAR footprint patch, an optical image patch, and an optical footprint patch. Figure 5 shows examples of the dataset on six buildings, of which three are intact and three are damaged.
@@ III Experimental Results
We benchmark an image classification task of two classes: damaged buildings and intact buildings. As the two classes are highly imbalanced between damaged and intact buildings, it can also be considered as an anomaly detection task, i.e., detecting damaged buildings in the whole dataset.
@@ III-A Baseline Approaches
Four methods are introduced to evaluate the performance and to establish a benchmark: support vector machine (SVM), random forests (RF), a 3-layer convolutional neural networks (CNN), and ResNet-18 [16].
SVM and RF are selected for their good performance in many applications, including classifying collapsed and standing buildings from post-event SAR imagery as reported in [11]. To extract features as the input, we follow the experimental setup proposed in [11], which employs four first-order statistics, i.e., mean, variance, skewness, and kurtosis, and eight second-order image statistical measures, i.e., mean, variance, homogeneity, contrast, dissimilarity, entropy, second moment, and correlation. For a detailed explanation, the readers are referred to [11].
A simple 3-layer CNN and a ResNet-18 are selected as the two deep learning methods. The simple CNN consists of 3 convolution-ReLU-maxpool blocks, followed by average pooling, a linear layer with dropout, and a final classification layer. The ResNet-18 follows the standard design in [16], consisting of 18 convolutional layers with residual connections. Similar to the simple CNN, we also add a linear layer with dropout before the final classification layer.
@@ III-B Evaluation Metrics
where $P$ and $R$ denote the precision and recall, respectively, and $tp$, $fp$, $tn$, $fn$ represent true positives, false positives, true negatives, and false negatives for buildings, respectively.
@@ III-C Implementation Details
We preprocess the images by removing 2% pixel outliers and normalizing the pixel values to the range [0,1]. For simple CNN, we randomly initialize the model; for ResNet-18, we use ImageNet pretrained weights. We use random resized crop and random horizontal and vertical flip as data augmentations. To deal with the significant class imbalance, we give a bigger weight to damaged buildings and a smaller weight to intact buildings during data sampling. We optimize binary cross entropy loss with AdamW optimizer for 30 epochs. The learning rate follows a cosine-decay schedule starting from 0.0001. Batch size is set to 32.
For the statistic features used by the two machine learning methods, we calculated the features of image patches based on the implementation of or PyFeats [^5] https://github.com/giakou4/pyfeats. For calculating the evaluation metrics, we use the implementation in scikit-learn [^6] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.roc_auc_score.html.
@@ III-D Performance Comparison
Table I illustrates variations in performance across different models on the dataset. For the SAR image, the SVM model shows the lowest precision but excels in recall (0.495), indicating it detects a high proportion of true positives but also has many false positives. RF outperforms other models in precision but has a lower recall. While having moderate precision and recall, CNN produces a relatively high $F_{1}$ score and the highest AUROC of 0.739 among SAR models, indicating its proficiency in handling complex features directly from raw SAR data. ResNet-18 underperforms CNN regarding all four metrics despite its higher complexity.
For the optical image, ResNet-18 stands out with the highest precision, recall, $F_{1}$ score, and AUROC. It outperforms other models across all metrics. The other deep learning model, CNN, demonstrates a balance between precision and recall (0.432), resulting in a high $F_{1}$ score and an impressive AUROC of 0.853. RF attains high precision at 0.611 but a relatively lower recall rate. In contrast, SVM yields less favorable outcomes compared to other models. Nevertheless, it is noteworthy that SVM’s AUROC of 0.723 surpasses that of three other models in the SAR image category.
Comparing SAR and optical images, we can see that SAR images appear to be more challenging for all models, with generally lower performance than optical images. For the two deep learning models, we observe that ResNet-18 excels in the optical images; however, CNN outperforms ResNet-18 regarding SAR images. The simple CNN’s better performance on SAR data might be attributed to several factors. SAR data’s unique characteristics may favor simpler models that efficiently capture essential patterns. The risk of overfitting is reduced with simpler models, which can lead to better generalization, particularly in cases of limited training data. In contrast, the complex deep neural networks may not be optimized for SAR-specific traits.
TABLE I: Benchmark results on the dataset: 5-fold mean (std). The highest values of different metrics are highlighted in bold.
image model Precision Recall $F_{1}$ AUROC
SAR SVM 0.096 (0.021) 0.495 (0.202) 0.154 (0.037) 0.653 (0.068)
RF 0.223 (0.090) 0.344 (0.208) 0.231 (0.071) 0.670 (0.079)
CNN 0.167 (0.028) 0.371 (0.073) 0.228 (0.032) 0.739 (0.026)
ResNet-18 0.155 (0.067) 0.276 (0.080) 0.184(0.050) 0.653 (0.043)
Optical SVM 0.163 (0.044) 0.352 (0.148) 0.207 (0.057) 0.723 (0.047)
RF 0.611 (0.215) 0.354 (0.098) 0.421 (0.106) 0.810 (0.062)
CNN 0.489 (0.154) 0.432 (0.078) 0.449 (0.096) 0.853 (0.045)
ResNet-18 0.666 (0.121) 0.539 (0.147) 0.581 (0.100) 0.938 (0.022)
@@ IV Conclusion
Detecting earthquake-damaged buildings in post-event satellite imagery is essential yet challenging. This study introduced a dataset designed to address the issue of data scarcity, a crucial factor in fostering the development of robust algorithms for this purpose. The dataset combines post-event SAR and optical satellite images with labels of damaged and intact buildings, and the problem is formulated as an image classification task. We provided a benchmark on both modalities with four baseline methods. Results show that detecting damages from post-event SAR images is valuable and possible but more challenging than optical images. Such findings call for further research on improved methods, in particular on SAR images.
@@
We refrained from data fusion as optical and SAR data are typically not simultaneously acquired. The dataset primarily aims to assess individual post-event SAR image performance and compare it with optical images. Nevertheless, the fusion of multi-modal data for post-earthquake building damage assessment may potentially enhance the overall effectiveness, offering a more comprehensive analysis of the damaged buildings when the data are available.
Acknowledgment
The authors would like to thank Capella Space (https://www.capellaspace.com/) and Maxar Technologies (https://www.maxar.com/) for providing satellite imagery and OpenStreetMap and Humanitarian OpenStreetMap Team (https://www.hotosm.org/) for providing building footprints and annotations of destroyed buildings.

### 05 Apr 2024 v2

@@ Abstract
This is the preprint version. Quick and automated earthquake-damaged building detection from post-event satellite imagery is crucial, yet it is challenging due to the scarcity of training data required for developing robust algorithms. This letter presents the first dataset dedicated to detecting earthquake-damaged buildings from post-event very high resolution (VHR) Synthetic Aperture Radar (SAR) and optical imagery. Utilizing open satellite imagery and annotations acquired after the 2023 Turkey–Syria earthquakes, we deliver a dataset of co-registered building footprints and satellite image patches of both SAR and optical data, encompassing more than four thousand buildings. The task of damaged building detection is formulated as a binary image classification problem, that can also be treated as an anomaly detection problem due to extreme class imbalance. We provide baseline methods and results to serve as references for comparison. Researchers can utilize this dataset to expedite algorithm development, facilitating the rapid detection of damaged buildings in response to future events. The dataset and codes together with detailed explanations and visualization will be made publicly available at https://github.com/ya0-sun/PostEQ-SARopt-BuildingDamage.
@@ I Introduction
Earthquakes can result in substantial structural and infrastructural damage, often with significant socioeconomic consequences. After an event, fast and accurate detection of earthquake-damaged buildings in remote sensing imagery is of great importance. Remote sensing technologies can effectively improve the efficiency of disaster management and have been employed to estimate the extent of earthquake damage to buildings [1, 2, 3]. Very high resolution (VHR) optical images are easier to interpret, making them a preferred choice for many studies. However, acquisition of cloud-free optical images depends on weather conditions and often needs to wait. In contrast, SAR imagery is particularly suitable for rapid disaster response scenarios as it can be acquired regardless of cloud coverage and sun illumination conditions. Moreover, the enhanced resolution of contemporary SAR satellite images enables the extraction of information at the individual building level [4, 5, 6], comparable to VHR optical data.
Over the past years, many researchers have developed algorithms utilizing SAR data to detect earthquake-damaged buildings. Most existing works detect changes using both pre- and post-earthquake SAR images to acquire building damage information [7, 8]; however, pre-event high-resolution SAR imagery is generally unavailable in most locations. A few works simulate a SAR image using building shapes extracted from a pre-event optical image and acquisition parameters of a post-event SAR image and detect earthquake-damaged buildings by comparing the simulated and real SAR images  [9]. Utilizing only a single post-event SAR image, some researchers detect damaged buildings by analyzing signatures of destroyed buildings in high-resolution SAR data [10, 11, 12]. However, the study areas often comprise a limited number of isolated buildings that do not depict the typical conditions in densely populated urban areas, where partially occluded buildings and geometric distortions in SAR images, i.e., foreshortening, layover, and shadowing, commonly exist. Therefore, the question of whether a single VHR SAR image acquired after an event can effectively identify damaged buildings remains to be addressed. In this regard, benchmark datasets play a pivotal role in the development and comparative assessment of diverse methodologies aimed at addressing the following questions: To what extent can a single VHR SAR image, acquired post-event, allow to identify damaged buildings and with which accuracy? Additionally, how do the outcomes derived from a single post-event SAR image compare with those of an optical image?
@@ II-A Study Area and Data Sources
The study area is chosen in the city of Islahiye, located in southeastern Turkey near the northwestern border of Syria. On February 6, 2023, a magnitude 7.8 earthquake struck Kahramanmaras, Turky, followed by a 7.5 magnitude aftershock nine hours later. The earthquakes inflicted widespread destruction, leading to significant damage to buildings, injuries, and loss of life. Islahiye was one of the most affected areas.
@@ II-B2 Registering corresponding features
Since the SAR image, i.e., the GEO product, has been terrain corrected, we neglect global registration and perform subarea and polygon registration. Afterward, manual validation and editing with expert knowledge are conducted to ensure precise registration, relying on identifying double bounce lines and the assumption of no abrupt terrain changes for editing buildings with unclear double bounce lines [13]. For individual buildings, we estimate a maximum registration error of 5-6 pixels (approximately 2 m) that is sufficient for locating individual buildings. Figure 2 shows building polygons on the SAR image in exemplary areas before and after registration. As can be seen, the registration procedure effectively aligns the two data sets. Note that buildings in the mountainous area on the west side of the city (e.g., in or near the red box in Figure 1) are excluded due to unclear signals in the SAR image, posing challenges for registration verification.
@@ II-C Patch generation
Among the 4029 buildings, 169 are damaged, and the remaining 3860 buildings are intact. Consequently, the dataset contains 169 damaged buildings and 3860 intact buildings, and each of them has four patches: a SAR image patch, a SAR footprint patch, an optical image patch, and an optical footprint patch. Figure 5 shows examples of the dataset on six buildings, of which three are intact and three are damaged.
TABLE I: Benchmark results on the dataset: 5-fold mean(std). The highest values of different metrics are highlighted in bold for each set of utilized data. Early Fusion is noted as EF, while Late Fusion is noted as LF.
Image SAR optical SAR+optical
model SVM RF CNN ResNet18 ResNet18 SVM RF CNN ResNet18 ResNet18 ResNet18
(EF) (EF+ImageNet) (LF+SAR-HUB) (EF) (EF+ImageNet) (LF+ImageNet) (LF+SAR-HUB+ImageNet)
Precision 0.096(0.021) 0.223(0.090) 0.167(0.028) 0.155(0.067) 0.237(0.116) 0.163(0.044) 0.611(0.215) 0.489(0.154) 0.666(0.121) 0.622(0.213) 0.746(0.062)
Recall 0.495(0.202) 0.344(0.208) 0.371(0.073) 0.276(0.080) 0.427(0.104) 0.352(0.148) 0.354(0.098) 0.432(0.078) 0.539(0.147) 0.646(0.143) 0.615(0.117)
$F_{1}$ 0.154(0.037) 0.231(0.071) 0.228(0.032) 0.184(0.050) 0.282(0.077) 0.207(0.057) 0.421(0.106) 0.449(0.096) 0.581(0.100) 0.605(0.120) 0.670(0.072)
AUROC 0.653(0.068) 0.670(0.079) 0.739(0.026) 0.653(0.043) 0.769(0.032) 0.723(0.047) 0.810(0.062) 0.853(0.045) 0.938(0.022) 0.941(0.026) 0.962(0.023)
@@ III Experimental Results
We benchmark an image classification task with two classes: damaged and intact buildings. Due to the significant class imbalance, it can also be viewed as an anomaly detection task, i.e., detecting damaged buildings within the entire dataset.
@@ III-A Baseline Approaches
Four models are introduced to establish a benchmark: support vector machine (SVM), random forest (RF), 3-layer convolutional neural network (CNN), and ResNet-18 [16].
SVM and RF are selected for their good performance in many applications, including classifying collapsed and standing buildings from post-event SAR imagery as reported in [11]. To extract features as the input, we mask out non-building pixels with footprints, and follow the setup proposed in [11], which employs four first-order statistics, i.e., mean, variance, skewness, and kurtosis, and eight second-order image statistical measures, i.e., mean, variance, homogeneity, contrast, dissimilarity, entropy, second moment, and correlation. For a detailed explanation, the readers are referred to [11].
A simple 3-layer CNN and a ResNet-18 are selected as the deep learning backbones. The simple CNN consists of 3 convolution-ReLU-maxpool blocks, followed by average pooling, a linear layer with dropout, and a final classification layer. We stack image and building footprint as the input. The ResNet-18 follows the standard design in [16], with which we benchmark both early and late fusion results of images and building footprints. For late fusion, images and building footprints are encoded by separate encoders, and the feature vectors are concatenated together to a following linear layer with dropout and a final classification layer. Apart from single-modal results, we also conduct a late fusion experiment with both SAR and optical data as a multimodal reference.
@@ III-B Evaluation Metrics
where $P$ and $R$ denote the precision and recall, and $tp$, $fp$, $tn$, $fn$ represent true positives, false positives, true negatives, and false negatives for buildings, respectively. The thresholds for positive/negative are determined by best $F_{1}$ scores.
@@ III-C Implementation Details
We preprocess the images by removing 2% pixel outliers and normalizing the pixel values to the range [0,1]. For simple CNN, we randomly initialize the model; for ResNet-18 early fusion, we use pure ImageNet weights; for ResNet-18 late fusion, we use ImageNet weights for optical images and footprints, and SAR-HUB [17] weights for SAR images. We use random resized crop and random horizontal and vertical flip as data augmentations. To deal with the significant class imbalance, we give a bigger weight to damaged buildings and a smaller weight to intact buildings during data sampling. We optimize binary cross entropy loss with AdamW optimizer for 30 epochs. The learning rate follows a cosine-decay schedule starting from 0.0001. Batch size is set to 32.
For the statistic features used by the two machine learning methods, we calculated the features of image patches based on the implementation of PyFeats [^5] https://github.com/giakou4/pyfeats. For calculating the evaluation metrics, we use the implementation in scikit-learn [^6] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.roc_auc_score.html.
@@ III-D Performance Comparison
Table I illustrates variations in performance across different models on the dataset. In general, deep neural networks outperform SVM and RF in both SAR and optical scenarios. For the SAR image, ResNet-18 underperforms simple CNN regarding all four metrics with early fusion and ImageNet weights, indicating the optimization challenge of complex SAR data. This issue is resolved with proper weight initialization from SAR-HUB (around 10% improvement in AUROC and $F_{1}$ scores), highlighting the importance of SAR pretrained models such as provided in [17].
For the optical image, ResNet-18 with late fusion stands out with the highest recall, $F_{1}$ score, and AUROC. It outperforms other models across these three metrics. ResNet-18 with early fusion gived best precision. The other deep learning model, CNN, demonstrates a balance between precision and recall, resulting in a high $F_{1}$ score and an impressive AUROC of 0.853. RF attains high precision at 0.611 but a relatively lower recall rate. In contrast, SVM yields less favorable outcomes compared to other models.
Comparing SAR and optical images, we can see that SAR images appear to be more challenging for all models, with generally lower performance than optical images. Nevertheless, good SAR models, e.g., ResNet-18 with SAR-HUB pretrained weights, outperform bad optical models such as SVM.
In addition, the fusion of SAR and optical imagery provides further improvement compared to each single modality, confirming the complementary information across different modalities.
@@ IV Conclusion
Detecting earthquake-damaged buildings in post-event satellite imagery is essential yet challenging. This study introduces a dataset designed to address the issue and to foster the development of robust algorithms. The dataset combines post-event SAR and optical satellite images with labels of damaged and intact buildings, and the problem is formulated as an image classification task. We provide a benchmark on both modalities with different baseline methods and a baseline of fusing optical and SAR data. Results show that detecting damage from post-event SAR images is valuable and possible but more challenging than optical images. Such findings call for further research on improved methods, in particular on SAR images. In addition, the performance gain through simple SAR-optical fusion verifies the potential in using multimodal data when they can be acquired following an event.